# Distributional Bellman Operators over Mean Embeddings

### 09 Dec 2023 v1

@@ 3.2 Sketch-DP at work
To provide more intuition for the Bellman sketch framework, we provide a walk-through of using Algorithm 1 to estimate the return distributions for the environment in Figure 2A; full details for replication are given in Appendix C. We take a feature map $\phi$ of the form given in Equation (8), taking $\kappa$ to be the sigmoid function, and $m=13$ anchors evenly spaced between $-4.5$ and $4.5$ (Figure 2B). The Bellman regression problem in Equation (5) is set with $\mu=\text{Uniform}([-4,4])$, based on the typical returns observed in the environment. The anchors and the choice of $\mu$ for regression are important but can be set following simple heuristics; see Sections B.2 and B.3. We then run the Sketch-DP algorithm with the initial estimates $U(x)$ set to $\phi(0)$ for all $x\in\mathcal{X}$.
We compare the estimates produced by Sketch-DP against ground-truth by estimating the true mean embeddings from a large number of Monte Carlo samples of the returns from each state. Figure 2C (top) shows illustrates the ground-truth embeddings for each state, and Figure 2D (top) compares these ground-truth embeddings with those computed by Sketch-DP as the algorithm progresses; by 30 iterations, the mean embeddings are very close to the ground-truth.
To aid interpretation of these results, we also include a comparison in which we “decode” the mean embeddings back into probability distributions (via an imputation strategy (Rowland et al., 2019)), and compare with the ground-truth return distributions, projected onto the anchor locations of the features (Rowland et al., 2018). Full details of the imputation strategy are in Section B.1. These results are shown in the bottom panels of Figure 2C & D. Initially, the imputed distributions of the Sketch-DP mean embedding estimates reflect the initialisation to the mean embedding of $\delta_{0}$, though as more iterations of Sketch-DP are applied, the imputed distributions become close to the ground-truth. This indicates that, in this example, not only does Sketch-DP compute accurate mean embeddings of the return, but that this embedding is rich enough to recover a lot of information regarding the return distributions themselves.
@@
ARefer to caption BRefer to caption
C                    DRefer to caption Refer to caption    Refer to caption    Refer to caption    Refer to caption    Refer to caption
Figure 2: A: State transitions and rewards in the environment. B: The feature functions $\phi$ for the sketch-DP. Dotted lines indicate anchors. The regression Equation (5) is performed under a densely spaced grid over the light region $[-4,4]$. C: The ground-truth mean embeddings under the sigmoid features in B, and the categorical projection of the ground-truth distribution onto the anchors of $\phi$ in B. D: The evolution of the estimated mean embeddings (bright blue lines) and imputed distributions (bright red lines) during Sketch-DP. The stems are the respective ground-truth from panel C.
@@ 4 Convergence analysis
[…]
Figure 3: The objects and structure used to analyse the Sketch-DP algorithm.
We analyse the Sketch-DP procedure described in Algorithm 1, which can be mathematically described in the following succinct manner. We let $U_{0}:\mathcal{X}\rightarrow\mathbb{R}^{m}$ denote the initial sketch value estimates, and then note from Algorithm 1 that the collection of estimates after each DP update form a sequence $(U_{k})_{k=0}^{\infty}$, with $U_{k+1}=\mathcal{T}^{\pi}_{\phi}U_{k}$. Our convergence analysis therefore focuses on the asymptotic behaviour of this sequence. We introduce the notation $\Phi:\mathscr{P}(\mathbb{R})\rightarrow\mathbb{R}^{m}$ for the sketch associated with the feature function $\phi$, so that $\Phi\mu=\mathbb{E}_{Z\sim\mu}[\phi(Z)]$, and define $\Phi$ for return-distribution functions by specifying for $\eta\in\mathscr{P}(\mathbb{R})^{\mathcal{X}}$ that $(\Phi\eta)(x)=\Phi(\eta(x))$. Ideally, we would like these iterates to approach $U^{\pi}:\mathcal{X}\rightarrow\mathbb{R}^{m}$, the sketch values of the true return distributions, given by $U^{\pi}(x)=\mathbb{E}^{\pi}_{x}[\phi(\sum_{t=0}^{\infty}\gamma^{t}R_{t})]$. As already described, typically this is not possible when the sketch $\Phi$ is not Bellman closed, and so we can only expect to approximate $U^{\pi}$. Mathematically, this is because in general we have $\Phi\mathcal{T}^{\pi}\not=\mathcal{T}^{\pi}_{\phi}\Phi$ when $\phi$ is not Bellman closed.
@@
$\max_{x\in\mathcal{X}}\|\Phi(\mathcal{T}^{\pi}\eta)(x)-(\mathcal{T}^{\pi}_{\phi}\Phi\eta)(x)\|\leq\sup_{g\in[G_{\text{min}},G_{\text{max}}]}\max_{r\in\mathcal{R}}\|\phi(r+\gamma g)-B_{r}\phi(g)\|\,.$ (9)
@@
  • (Bellman approximation bound.) For any $\eta\in\mathscr{P}([G_{\text{min}},G_{\text{max}}])^{\mathcal{X}}$,
    $\max_{x\in\mathcal{X}}\|\Phi(\mathcal{T}^{\pi}\eta)(x)-(\mathcal{T}^{\pi}_{\phi}\Phi\eta)(x)\|\leq\varepsilon_{\text{B}}\,.$
  • (Reconstruction error bound.) For any $\eta,\bar{\eta}\in\mathscr{P}([G_{\text{min}},G_{\text{max}}])^{\mathcal{X}}$ with sketches $U,\bar{U}$, we have $d(\eta,\bar{\eta})\leq\|U-\bar{U}\|_{\infty}+\varepsilon_{\text{R}}$.
  • (Embedding error bound.) For any $\eta^{\prime},\bar{\eta}^{\prime}\in\mathscr{P}([G_{\text{min}},G_{\text{max}}])^{\mathcal{X}}$ with sketches $U^{\prime},\bar{U}^{\prime}$, we have $\|U^{\prime}-\bar{U}^{\prime}\|_{\infty}\leq d(\eta^{\prime},\bar{\eta}^{\prime})+\varepsilon_{\text{E}}$.
Then for any two return-distribution functions $\eta,\bar{\eta}\in\mathscr{P}([G_{\text{min}},G_{\text{max}}])^{\mathcal{X}}$ with sketches $U,\bar{U}$ satisfying $\|U-\bar{U}\|\leq\delta$, we have
$\|\Phi\mathcal{T}^{\pi}\eta-\mathcal{T}^{\pi}_{\phi}\bar{U}\|_{\infty}\leq\gamma^{c}(\delta+\varepsilon_{\text{R}})+\varepsilon_{\text{R}}+\varepsilon_{\text{E}}\,.$
A formal proof is given in Appendix A; Figure 3 (bottom) shows the intuition, propagating bounds through different intermediate stages of the analysis of the update. We now state the main error bound result, which combines the two earlier results.
@@
Suppose the assumptions of Proposition 4.2 hold, that $\mathcal{T}^{\pi}$ maps $\mathscr{P}([G_{\text{min}},G_{\text{max}}])^{\mathcal{X}}$ to itself, and suppose $\mathcal{T}^{\pi}_{\phi}$ maps $\{\Phi\nu:\nu\in\mathscr{P}([G_{\text{min}},G_{\text{max}}])^{\mathcal{X}}\}$ to itself. Then for a sequence of sketches $(U_{k})_{k=0}^{\infty}$ defined iteratively via $U_{k+1}=\mathcal{T}^{\pi}_{\phi}U_{k}$, we have
$\limsup_{k\rightarrow\infty}\|U_{k}-U^{\pi}\|\leq\frac{1}{1-\gamma^{c}}(\gamma^{c}\varepsilon_{\text{R}}+\varepsilon_{\text{B}}+\varepsilon_{\text{E}})\,.$
@@ 4.1 Concrete example
$\limsup_{k\rightarrow\infty}\|U_{k}-U^{\pi}\|_{\infty}\leq\frac{(G_{\text{max}}-G_{\text{min}})(3+2\gamma)}{(1-\gamma)m}\,.$
@@ 5 Experiments
We first conduct a broad empirical investigation into the effects of three key factors in Equation (8): the base feature $\kappa$, the number of features $m$, and the slope $s$, using three tabular MRPs (details in Section C.1, extended results in Section D.1). As in the example in Section 3.2, we compare the mean embeddings estimated by Sketch-DP with ground-truth mean embeddings, reporting their squared $L^{2}$ distance (mean embedding squared error), and also compare the Cramér distance $\max_{x\in\mathcal{X}}\ell_{2}^{2}(\hat{\eta}(x),\eta^{\pi}(x))$ (see e.g. Rowland et al. (2018)) between the distribution $\hat{\eta}(x)$ imputed from the Sketch-DP estimate, and the ground-truth return distribution $\eta^{\pi}(x)$. To aid interpretation of the Cramér distance results, we also report the Cramér distance between the ground truth $\eta^{\pi}(x)$ and two baselines. First, the Dirac delta $\delta_{V^{\pi}(X)}$ at the mean return; we expect Sketch-DP to outperform this naïve baseline by better capturing properties of the return distribution beyond the mean. Second, the return distribution estimate computed by categorical DP (Rowland et al., 2018; Bellemare et al., 2023), a well-understood approach to distrbutional RL based on categorical distributions.
@@
The results for sweeps over feature count $m$ and slope $s$ are shown in Figure 4. By sweeping over $m$, we see that the estimated mean embedding goes towards the ground-truth as we use more features. Further, the Cramér distance also decreases as $m$ increases, suggesting that the distribution represented also approaches the ground-truth. To highlight differences between various Sketch-DP algorithms, we also compute the excess Cramér: the Cramér distance $\max_{x\in\mathcal{X}}\ell_{2}^{2}(\hat{\eta}(x),\eta^{\pi}(x))$ as above, minus the corresponding distance between the categorical projection of $\eta^{\pi}$ (c.f. the red stems in Figure 2) and $\eta^{\pi}$ itself. All distributional methods perform well on these tasks, and significantly outperform the Dirac estimator in stochastic environments; we note that all methods have tunable hyperparameters (bin locations for CDRL, feature parameters for Sketch-DP), which should inform the interpretation of these results, and in particular direct comparison between methods. The results of the sweep on the slope parameter $s$ show different trends depending on the metric. For smoother $\phi$, generally we can obtain smaller error on the mean embeddings, but the Cramér distances are only small for intermediate range of slope values. This result is expected: when the features are too smooth or too sharp, there exists regions within the return range where the feature values do not vary meaningfully. This results in a more lossy encoding of the return distribution, indicating the importance of tuning the slope parameter of the translation family (Equation (8)).
Figure 4: Results of running Algorithm 1 on tabular environments.
5.1 Deep reinforcement learning
We also verify that the Bellman sketch framework is robust enough to apply in combination with deep reinforcement learning. To do so, we aim to learn neural-network predictions $U_{\theta}(x,a)$ of sketch values for each state-action pair $(x,a)$ in the environment. To be able to define greedy policy improvements based on estimated sketch values, we precompute value-readout coefficients $\beta\in\mathbb{R}^{m}$ by solving $\operatorname*{arg\,min}_{\beta}\mathbb{E}_{G\sim\mu}[(G-\langle\beta,\phi(G)\rangle)^{2}]\,,$ so that we can predict expected returns from the sketch value as $\langle\beta,U_{\theta}(x,a)\rangle$. This allows us to define a greedy policy, and therefore a Q-learning-style update rule, which given an observed transition $(x,a,r,x^{\prime})$, first computes $a^{\prime}=\operatorname*{arg\,max}_{\tilde{a}}\langle\beta,U_{\bar{\theta}}(x^{\prime},\tilde{a})\rangle$, and then the gradient: $\nabla_{\theta}\|U_{\theta}(x,a)-B_{r}U_{\bar{\theta}}(x^{\prime},a^{\prime})\|^{2}_{2}\,,$ where $\bar{\theta}$ are the target network parameters. In our experiments, we parametrise $U_{\theta}$ according to the architecture of QR-DQN (Dabney et al., 2018b), so that the $m$ outputs of the network predict the values of the $m$ coordinates of the corresponding sketch value. We use the sigmoid function as the base feature $\kappa$. Full experimental details for replication are in Appendix C.2; further results are in Section D.2.
@@ 5.1 Deep reinforcement learning
Figure 5 shows the mean and median human-normalised performance on the Atari suite of environments (Bellemare et al., 2013) across 200M training frames, and includes comparisons against DQN (Mnih et al., 2015), as well as the distributional agents C51 (Bellemare et al., 2017), QR-DQN (Dabney et al., 2018b), and IQN (Dabney et al., 2018a). Sketch-DQN attains higher performance on both metrics relative to the comparator agents C51 and QR-DQN, and approaches the performance of IQN, which uses a more complex prediction network to make non-parametric predictions of the quantile function of the return. These results indicate that the sketch framework can be reliably applied to deep RL, and we believe further investigation of the combination of this framework and deep RL agents is a promising direction for future work.
@@ 6 Related work
Typical approaches to distributional RL focus on learning approximate distributions directly (see, e.g., Bellemare et al. (2017); Dabney et al. (2018b); Yang et al. (2019); Nguyen-Tang et al. (2021); Wu et al. (2023)). Much prior work has considered statistical functionals of the random return, at varying levels of generality with regard to the underlying Markov decision process model. See for example Mandl (1971); Farahmand (2019) for work on characteristic functions, Chung & Sobel (1987) for the Laplace transform, Tamar et al. (2013; 2016) for variance, and Sobel (1982) for higher moments. Our use of finite-dimensional mean embeddings is inspired by distributed distributional codes (DDCs) from theoretical neuroscience (Sahani & Dayan, 2003; Vértes & Sahani, 2018; Wenliang & Sahani, 2019), which can be regarded as neural activities encoding return distributions. DDCs were previously used to model transition dynamics and successor features in partially observable MDPs (Vértes & Sahani, 2019). Tano et al. (2020) consider applying non-linearities to rewards themselves, rather than the return, and learning with a variety of discount factors, to encode the distribution of rewards at each timestep. The sketches in this paper are in fact mean embeddings into finite-dimensional reproducing kernel Hilbert spaces (RKHSs; the kernel corresponding to the feature function $\phi$ is $K(z,z^{\prime})=\langle\phi(z),\phi(z^{\prime})\rangle$). Kernel mean embeddings have previously been used in RL for representing state-transition distributions (Grünewälder et al., 2012; Boots et al., 2013; Lever et al., 2016; Chowdhury & Oliveira, 2023), and maximum mean discrepancies in RKHSs (Gretton et al., 2012) have been used to define losses in distributional RL by Nguyen-Tang et al. (2021).

### 04 Mar 2024 v3

@@ 3.2 Sketch-DP at work
To give more intuition for the Bellman sketch framework, we provide a walk-through of using Algorithm 1 to estimate the return distributions for the environment in Figure 2A. We take a sinusoidal feature map $\phi$ that consists of $m=13$ harmonics over the range $[-4.5,4.5]$ (see Section D.2 for an example using feature map of the form Equation (8)). The Bellman regression problem in Equation (5) is set with $\mu=\text{Uniform}([-4,4])$, based on the typical returns observed in the environment. We then run the Sketch-DP algorithm with the initial estimates $U(x)$ set to $\phi(0)$ for all $x\in\mathcal{X}$.
To visualise how the estimated mean embeddings evolve over iterations, we project them onto their first two principal components in Figure 2C, which explain $>70\%$ variance. To approximate the ground truth return distributions, we collected a large number of Monte Carlo samples from the MRP Figure 2D; see Section C.1 for details. We then estimate the ground-truth mean embeddings and project them on to the principal subspace in Figure 2C as crosses. The Sketch-DP estimates converge to close proximity of the ground-truth. The distinctive update pattern stems from the fact that all paths between rewarding states have length 3. The mean embeddings of states 1 and 2 are closer to state 4 due to more frequent transitions from state 2 to 4.
To aid interpretation of these results, we also include a comparison in which we “decode” the mean embeddings at selected iterations back into probability distributions (via an imputation strategy (Rowland et al., 2019)), and compare with the ground-truth return distributions projected onto the anchor locations of the features (Rowland et al., 2018), as shown in Figure 2D. Full details of the imputation strategy are in Section B.1. These decoded distributions are shown in Figure 2E. Initially, the imputed distributions of the Sketch-DP mean embedding estimates reflect the initialisation to the mean embedding of $\delta_{0}$. As more iterations of Sketch-DP are applied, the imputed distributions of the evolving mean embedding estimates become close to the projected ground-truth. This indicates that, in this example, not only does Sketch-DP compute accurate mean embeddings of the return, but that this embedding is rich enough to recover a lot of information regarding the return distributions themselves.
@@ 4 Convergence analysis
We analyse the Sketch-DP procedure described in Algorithm 1, with a novel error analysis approach that can be mathematically described in the following succinct manner. We let $U_{0}:\mathcal{X}\rightarrow\mathbb{R}^{m}$ denote the initial sketch value estimates, and then note from Algorithm 1 that the collection of estimates after each DP update form a sequence $(U_{k})_{k=0}^{\infty}$, with $U_{k+1}=\mathcal{T}^{\pi}_{\phi}U_{k}$. Our convergence analysis therefore focuses on the asymptotic behaviour of this sequence. We introduce the notation $\Phi:\mathscr{P}(\mathbb{R})\rightarrow\mathbb{R}^{m}$ for the sketch associated with the feature function $\phi$, so that $\Phi\mu=\mathbb{E}_{Z\sim\mu}[\phi(Z)]$, and define $\Phi$ for return-distribution functions (RDFs) by specifying for $\eta\in\mathscr{P}(\mathbb{R})^{\mathcal{X}}$ that $(\Phi\eta)(x)=\Phi(\eta(x))$. Ideally, we would like these iterates to approach $U^{\pi}:\mathcal{X}\rightarrow\mathbb{R}^{m}$, the sketch values of the true return distributions, given by $U^{\pi}(x)=\mathbb{E}^{\pi}_{x}[\phi(\sum_{t=0}^{\infty}\gamma^{t}R_{t})]$. As already described, typically this is not possible when the sketch $\Phi$ is not Bellman closed, and so we can only expect to approximate $U^{\pi}$. Mathematically, this is because in general $\Phi\mathcal{T}^{\pi}\not=\mathcal{T}^{\pi}_{\phi}\Phi$ when $\phi$ is not Bellman closed.
@@
$\max_{x\in\mathcal{X}}\|$ $\Phi(\mathcal{T}^{\pi}\eta)(x)-(\mathcal{T}^{\pi}_{\phi}\Phi\eta)(x)\|$ (9)
[…]
Figure 3: The objects and structure used to analyse the Sketch-DP.
@@
[…]
A formal proof is given in Appendix A; Figure 3 (bottom) shows the intuition, propagating bounds through different intermediate stages of the analysis of the update. The main error bound result combines the two earlier results.
@@
[…]
@@ 4.1 Concrete example
[…]
Figure 4: Results of running Sketch-DP (Algorithm 1) on tabular environments.
@@ 5 Experiments
We first test quality of return distribution predictions by the sketch algorithms, investigating the effects of three key factors in Equation (8): the base feature $\kappa$, the number of features $m$, and the slope $s$, using three tabular MRPs (details in Section C.1, extended results in Section D.3). In Random chain, the transitions are random, rewards are deterministic; in Directed chain (DC), both transitions and rewards are deterministic; and in DC+Gaussian R, the transitions are deterministic and the rewards are Gaussian. We compare the mean embeddings estimated by Sketch-DP with ground-truth mean embeddings, reporting their squared $L^{2}$ distance (mean embedding squared error), and also compare the Cramér distance $\max_{x\in\mathcal{X}}\ell_{2}^{2}(\hat{\eta}(x),\eta^{\pi}(x))$ (see e.g. Rowland et al. (2018)) between the distribution $\hat{\eta}(x)$ imputed from the Sketch-DP estimate against the ground-truth $\eta^{\pi}(x)$. To aid interpretation of the Cramér distance results, we also report the Cramér distance between the ground truth $\eta^{\pi}(x)$ and two baselines. First, the Dirac delta $\delta_{V^{\pi}(X)}$ at the mean return; we expect Sketch-DP to outperform this naïve baseline by better capturing properties of the return distribution beyond the mean. Second, the return distribution estimate computed by categorical DP (Rowland et al., 2018; Bellemare et al., 2023), a well-understood approach to distrbutional RL based on categorical distributions.
@@
The results for sweeps over feature count $m$ and slope $s$ are shown in Figure 4. By sweeping over $m$, we see that the estimated mean embedding goes towards the ground-truth as we use more features. Further, the Cramér distance also decreases as $m$ increases, suggesting that the distribution represented also approaches the ground-truth. To highlight differences between various Sketch-DP algorithms, we also compute the excess Cramér: the Cramér distance as above, minus the corresponding distance between the categorical projection of $\eta^{\pi}$ and $\eta^{\pi}$ itself. All distributional methods perform well on these tasks, and significantly outperform the Dirac estimator in stochastic environments; we note that all methods have tunable hyperparameters (bin locations for CDRL, feature parameters for Sketch-DP), which should inform direct comparison between methods. The results of the sweep on the slope parameter $s$ show different trends depending on the metric. For smoother $\phi$, generally we can obtain smaller error on the mean embeddings, but the Cramér distances are only small for intermediate range of slope values. This result is expected: when the features are too smooth or too sharp, there exists regions within the return range where the feature values do not vary meaningfully. This results in a more lossy encoding of the return distribution, indicating the importance of tuning the slope parameter of the translation family (Equation (8)).
We include additional experiments in Section D.4 showing that Sketch-DP outperforms and is substantially faster in wallclock run-time than SFDP based on imputation strategies (Rowland et al., 2019; Bellemare et al., 2023). The faster speed of Sketch-DP is because its updates involve simple linear-algebraic operations, as opposed to the more involved DP update, using imputation strategies, in SFDP.
@@
5.1 Deep reinforcement learning
The primary motivation of our work has been to develop principled novel approaches to distributional RL based on mean embeddings. Here, we also verify that the Bellman sketch framework is robust enough to apply in combination with deep reinforcement learning. We train neural-network predictions $U_{\theta}(x,a)$ of sketch values for each state-action pair $(x,a)$. To be able to define greedy policy improvements based on estimated sketch values, we precompute value-readout coefficients $\beta\in\mathbb{R}^{m}$ by solving $\operatorname*{arg\,min}_{\beta}\mathbb{E}_{G\sim\mu}[(G-\langle\beta,\phi(G)\rangle)^{2}]\,,$ so that we can predict expected returns from the sketch value as $\langle\beta,U_{\theta}(x,a)\rangle$. This allows us to define a Q-learning-style update rule: given a transition $(x,a,r,x^{\prime})$, first compute $a^{\prime}=\operatorname*{arg\,max}_{\tilde{a}}\langle\beta,U_{\bar{\theta}}(x^{\prime},\tilde{a})\rangle$, and then the gradient: $\nabla_{\theta}\|U_{\theta}(x,a)-B_{r}U_{\bar{\theta}}(x^{\prime},a^{\prime})\|^{2}_{2}\,,$ where $\bar{\theta}$ are the target network parameters. In our experiments, we parametrise $U_{\theta}$ according to the architecture of QR-DQN (Dabney et al., 2018b), so that the $m$ outputs of the network predict the values of the $m$ coordinates of the corresponding sketch value. We use the sigmoid function as the base feature $\kappa$. Full experimental details for replication are in Appendix C.2; further results are in Section D.5.
Figure 5 shows the mean and median human-normalised performance on the Atari suite of environments (Bellemare et al., 2013) across 200M training frames, and includes comparisons against DQN (Mnih et al., 2015), as well as the distributional agents C51 (Bellemare et al., 2017), QR-DQN (Dabney et al., 2018b), and IQN (Dabney et al., 2018a). Sketch-DQN attains higher performance on both metrics relative to the comparator agents C51 and QR-DQN, and approaches the performance of IQN, which uses a more complex prediction network to make non-parametric predictions of the quantile function of the return. In addition, Sketch-DQN runs faster than QR-DQN and IQN; see Section D.6. These results indicate that the sketch framework can be reliably applied to deep RL. Code is available here.
@@ 6 Related work
Typical approaches to distributional RL focus on learning approximate distributions directly (see, e.g., Bellemare et al. (2017); Dabney et al. (2018b); Yang et al. (2019); Nguyen-Tang et al. (2021); Wu et al. (2023)). Much prior work has considered statistical functionals of the random return, at varying levels of generality with regard to the underlying Markov decision process model. See for example Mandl (1971); Farahmand (2019) for work on characteristic functions, Chung & Sobel (1987) for the Laplace transform, (Tamar et al., 2013, 2016) for variance, and Sobel (1982) for higher moments. Our use of finite-dimensional mean embeddings is inspired by distributed distributional codes (DDCs) from theoretical neuroscience (Sahani & Dayan, 2003; Vértes & Sahani, 2018; Wenliang & Sahani, 2019), which can be regarded as neural activities encoding return distributions. DDCs were previously used to model transition dynamics and successor features in partially observable MDPs (Vértes & Sahani, 2019). Tano et al. (2020) consider applying non-linearities to rewards themselves, rather than the return, and learning with a variety of discount factors, to encode the distribution of rewards at each timestep. The sketches in this paper are in fact mean embeddings into finite-dimensional reproducing kernel Hilbert spaces (RKHSs; the kernel corresponding to the feature function $\phi$ is $K(z,z^{\prime})=\langle\phi(z),\phi(z^{\prime})\rangle$). Kernel mean embeddings have been used in RL for representing state-transition distributions (Grünewälder et al., 2012; Boots et al., 2013; Lever et al., 2016; Chowdhury & Oliveira, 2023), and maximum mean discrepancies (MMDs) (Gretton et al., 2012) have been used to define losses in distributional RL by Nguyen-Tang et al. (2021). Nguyen-Tang et al. (2021) combine an MMD loss and distributional bootstrapping to define an incremental learning algorithm, but it does not naturally lead to a DP formulation, and its convergence is not analysed. In contrast, our sketch framework not only generalises their approach to define both DP and TD algorithms but also allows rigorous analysis of DP convergence as presented in Section 4. We elaborate the comparisons to MMDRL and a few other distributional RL methods in Section B.7.